OpenStudy (anonymous):

A local hamburger shop sold a combined total of 566 hamburgers and cheeseburgers on Saturday. There were 66 more cheeseburgers sold than hamburgers. How many hamburgers were sold on Saturday?

OpenStudy (anonymous):

cornzyblack.. i dont understand word problems like this..

OpenStudy (cornzyblack):

Ok ama try, first lets understand that the shop consisted of Cheeseburgers, and Hamburgers So let Hamburgers = H Cheeseburgers = C So 566 = C + H -------- (equation 1) Now since 66 more Cheeseburgers were sold more than Hamburgers, it means that, C = H + 66 so Substitute into eqn 1 566 = H + 66 + H 566 - 66 = 2H 500 = 2H H = 250 250 Hamburgers were sold on Saturday

OpenStudy (anonymous):

thankkkksssss

OpenStudy (cornzyblack):

welcome

OpenStudy (anonymous):

omg i just cant do these..

OpenStudy (anonymous):

A washer and a dryer cost 851 combined. The washer costs 99 less than the dryer. What is the cost of the dryer?

OpenStudy (anonymous):

ii gooottt it

OpenStudy (anonymous):

475

OpenStudy (anonymous):

i followed ur steps!!

OpenStudy (cornzyblack):

lol... it just actually involes you knowing what the figure consists of.. Like we know the shop sold 566 in saturday.. 566 what??? well 566 consisted of Hamburgers and Cheeseburgers So we just represent 566 = H + C Were H = Hamburgers C= Cheeseburgers and tardaaah